leila1234567:

Both Andrew and Karleigh recorded the distance they ran in x minutes on treadmills. Andrew A 2-column table with 2 rows. Column 1 is labeled Time in Minutes with entries 18, 24. Column 2 is labeled Distance in miles with entries 1.5, 2. Karleigh A 2-column table with 2 rows. Column 1 is labeled Time in Minutes with entries 30, 40. Column 2 is labeled Distance in miles with entries 3, 4. Andrew and Karleigh each run for 1 hour. Which statement explains who ran a greater distance? Andrew ran a greater distance. The slope of the line described by the data in his table increased at a rate of StartFraction 1 Over 10 EndFraction mile per minute compared to Karleigh’s StartFraction 1 Over 12 EndFraction mile per minute. Karleigh ran a greater distance. The slope of the line described by the data in her table increased at a rate of StartFraction 1 Over 10 EndFraction mile per minute compared to Andrew’s StartFraction 1 Over 12 EndFraction mile per minute. Karleigh ran a greater distance. The slope of the line described by the data in her table increased at a rate of StartFraction 1 Over 12 EndFraction mile per minute compared to Andrew’s StartFraction 1 Over 10 EndFraction mile per minute. Andrew ran a greater distance. The slope of the line described by the data in his table increased at a rate of StartFraction 1 Over 12 EndFraction mile per minute compared to Karleigh’s StartFraction 1 Over 10 EndFraction mile per minute.

leila1234567:

1 attachment
leila1234567:

1 attachment
giannawatson:

one sec

leila1234567:

kk

giannawatson:

so ignore the top look what the question is asking then i will tell you if your answer is wrong

leila1234567:

kk

giannawatson:

The one who ran at the greater Average Speed in 1 hour ran the greater Distance.

leila1234567:

so we have to do the slope equation?

leila1234567:

is it karleigh

giannawatson:

ex:Speed (S) is the ratio of the distance (D) covered to the time (t) taken. That is, S = D/t Suppose Andrew ran a distance D1 in 1 hour (3600 seconds) at a Speed, say S1, we have S1 = D1/t We can then say he ran a distance D1 = t × S1 = 3600S1 Similarly, let's say Karleigh ran a distance D2 = t × S2 = 3600S2 Let us compare these two, you will notice that the bigger number between S1 and S2 is going to determine the bigger number between D1 and D2. Let's choose random numbers for S1 and S2 for clarity, say S1 = 5, S2 = 10 D1 = 3600 × 5 = 18000 D2 = 3600 × 10 = 36000 This makes D2 bigger than D1. this is an example i found on the internet.

giannawatson:

hope that helps

leila1234567:

yep :3